OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10, 5/3, -5/18 ...

OpenStudy (anonymous):

Looks like \(\left\{a_n\right\}=\left\{(-1)^{n+1}\dfrac{60}{6^{n-1}}\right\}_{n=1}^\infty\). Find \(\displaystyle \lim_{n\to\infty}a_n\).

OpenStudy (anonymous):

what? @SithsAndGiggles I'm a bit confused

OpenStudy (anonymous):

The first thing I did was write the nth term. Another way of expressing the same sequence. You have alternating signs of successive terms, so you know there's a power of (-1) somewhere along the line, such as \((-1)^n\) or \((-1)^{n+1}\). Which of these you use depends on the sign of the first term. Your first term is positive 60. Starting your sequence with n=1, the only way to start with a positive number is if you have an even power of -1, so you use \((-1)^{n+1}\). (Plugging in n=1 gives you \((-1)^2=1\), thus a positive sign.) Now let's ignore the sign for now, since it's been taken care of. The first term is 60, then 60/6, then 60/36, and so on. In the denominator you have a power of 6 (\(6^n\)); each successive term is the previous term divided by 6. You start with \(\dfrac{60}{1}=\dfrac{60}{6^0}\). Since we're starting the sequence with n=1, we change the exponent on the 6 to n-1 to include this first term. So you're left with \[a_n=(-1)^{n+1}\frac{60}{6^{n-1}}\] Does that make sense?

OpenStudy (anonymous):

okay yes that makes sense! from that point how do I find the limit and decide if it converges or diverges

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

\[\lim_{n\to\infty}a_n=\lim_{n\to\infty}(-1)^{n+1}\frac{60}{6^{n-1}}\] Do you need help with the actual limit-finding?

OpenStudy (anonymous):

let me try it real quick

OpenStudy (anonymous):

oh wait yeah I need help finding it.. I'm not sure what to do with n

OpenStudy (anonymous):

\[\lim_{n\to\infty}(-1)^{n+1}\frac{60}{6^{n-1}}\] The \((-1)^{n+1}\) part will alternate between -1 and 1 as n becomes arbitrarily large, so think of the limit as \[\lim_{n\to\infty}\pm\frac{60}{6^{n-1}}\] The bottom will approach ∞, so...

OpenStudy (anonymous):

I am so lost!

OpenStudy (anonymous):

:( blehhh I'm going to die

OpenStudy (anonymous):

Back up a bit. According to the directions, you have to show whether the sequence is convergent, then provide the limit. Do you know how to show it's convergent?

OpenStudy (anonymous):

No I have no idea, I'm doing this class on my own and nothing covered it in my text it just went over geometric sequences :(

OpenStudy (anonymous):

That's actually exactly what you need! This sequence is geometric. What do you know about the convergence of a geo seq?

OpenStudy (anonymous):

I have to brb for about 15 minutes will u be on when i get back? I'm learning a lot from your explanation!

OpenStudy (anonymous):

Sure, I can wait.

OpenStudy (anonymous):

While I am waiting, I suppose I can write some more... \[\left\{a_n\right\}=\left\{(-1)^{n+1}\dfrac{60}{6^{n-1}}\right\}_{n=1}^\infty\] See why it's a geo seq? \[(-1)^{n+1}\dfrac{60}{6^{n-1}}=(-1)^{n+1}\frac{60\cdot6}{6^n}=360(-1)^{n+1}\left(\color{red}{\frac{1}{6}}\right)^n\] So you have a geo seq with \(r=\dfrac{1}{6}\). Does it converge or diverge? Once you answer this question, go back to finding the limit.